# First loosely coherent search for continuous gravitational wave sources with substellar companions in the Orion spur

### v2

@@ I Introduction
All-sky searches for continuous gravitational waves from isolated stars over extended observation periods are computationally demanding, as optimal search methods are often prohibitively expensive. Sub-optimal methods are typically employed, requiring trade-offs between search depth, breadth, and robustness. The problem becomes harder when the emitting object is in a binary system, because the orbital parameters now also need to be searched over. The result is that, with the same computational budget, an all-sky search for emission from a star in a binary system is less sensitive than an all-sky search from an isolated object over the same set of possible signal frequencies.
@@
We conducted a wide-band search of a portion of the sky, restricting the modulation depth to substellar companions. This is a physically interesting region that Falcon can search with a comparatively small amount of computational resources - less than 1.2 million CPU-thread hours, including follow-up analysis.
The follow-up analysis produced no outliers, which is unsurprising given that the searched sky area contains no hardware-injected simulated signals.
@@ II Search setup
Figure 2: Orbital parameters of known pulsars (circles), regions of orbital parameter space surveyed so far by continuous wave searches (dashed yellow lines) and by this search (shaded region). This figure is adapted from [4] .
@@ III Testing
Figure 3: Polarization-specific upper limits obtained for software injections. The red line at the bottom is a diagonal $y=x$.
As expected, the Falcon pipeline required modifications to adapt to the higher dimensional parameter spaces of the binary search. To validate its performance, we carried out several simulations where a signal was added to the O3 data and recovered with the new pipeline.
Figure 4: Recovery of software injections. The X axis shows software injection strain normalized to the worst-case upper limit obtained in absence of software injections. The Y axis shows the percentage of injections recovered. The vertical green line indicates when the software injection strain is equal to the upper limit obtained without software injection. The horizontal green line shows the 95% recovery level.
Figure 3 shows the results of an extensive validation study, where we computed polarization-specific upper limits [21] and compared them with the strain of the software injections. Polarisation-specific upper limits were obtained by the same procedure [22] as used in the actual analysis and contributing to the atlas. As we can see, the upper limits exceed the injected values. Worst-case and circularly polarized upper limits were verified as well.
The software injections were further used to verify follow-up performance. Figure 4 demonstrates that we recover 95% of injections above the worst-case upper limit level computed for the data without the injection.

### v3

@@ I Introduction
All-sky searches for continuous gravitational waves from isolated stars over extended observation periods are computationally demanding. As optimal search methods are prohibitively expensive, sub-optimal methods are employed, requiring trade-offs between search depth, breadth, and robustness. The problem becomes harder when the emitting object rather than being isolated is in a binary system, because the orbital parameters now also need to be searched over. The result is that, with the same computational budget, an all-sky search for emission from a star in a binary system is less sensitive than an all-sky search from an isolated object over the same set of possible signal frequencies.
@@
We conducted a wide-band search of a portion of the sky within the Orion Spur, restricting the modulation depth to substellar companions. This is a physically interesting region that Falcon can search with a comparatively small amount of computational resources - less than 1.2 million CPU-thread hours, including follow-up analysis.
The follow-up analysis produced no outliers. No outliers were expected from hardware injected simulated signals because none were present in the searched sky area.
@@ II Search setup
Figure 2: Orbital parameters of known pulsars (circles), regions of orbital parameter space surveyed so far by continuous wave searches (dashed yellow lines) and by this search (shaded region). The sensitivities of previous searches vary considerably, for example, the search convering the large rectangle was performed on pre-Advanced LIGO data. We also show contours of constant companion mass $m$ and the contours of companion orbital velocity $v$.
@@ III Testing
Figure 3: Polarization-specific upper limits obtained for test signals. The red line at the bottom is a diagonal $y=x$.
As expected, the Falcon pipeline required modifications to adapt to the higher dimensional parameter spaces of the binary search. To validate its performance, we carried out several simulations where test signals were added to the O3 data and recovered with the new pipeline.
Figure 4: Recovery of test signals. The X axis shows test signal strain normalized to the worst-case upper limit obtained in absence of test signals. The Y axis shows the percentage of injections recovered. The vertical green line indicates when the test signal strain is equal to the upper limit obtained without test signal. The horizontal green line shows the 95% recovery level.
Figure 3 shows the results of an extensive validation study, where we computed polarization-specific upper limits [21] and compared them with the strain of the added test signals. Polarisation-specific upper limits were obtained by the same procedure [22] as used in the actual analysis and contributing to the atlas. As we can see, the upper limits exceed the test-signal amplitudes. Worst-case and circularly polarized upper limits were verified as well.
The test signals were further used to verify follow-up performance. Figure 4 demonstrates that we recover 95% of test signals above the worst-case upper limit level computed for the data without the test signals.